### v1

@@ 5 Conclusion
In this paper, we focus on two of the variants of coflow scheduling with rational demand matrices, for which $O(1)$-approximations were not previously known: Namely, we present a 16-approximation algorithm for minimizing the average completion time under indirect fractional matching routings. Secondly, we address both the makespan and the average completion time objectives under indirect integral routing, and provide a full characterization of the worst-case runtimes under both objectives for this regime. While this falls short of providing $O(1)$-approximations for indirect integral routing, our proposed algorithm provides $O(\log n)$-approximations for the two objectives, in addition to being worst-case optimal; we also believe that the worst-case analysis will be helpful in determining whether a $O(1)$-approximation for indirect integral routing is possible. If it is, then we will have shown, together with the results in this paper, that constant-approximation algorithms exist for all possible coflow scheduling variants in Table˜1.

### v2

@@ 5 Conclusion
In this paper, we focus on two of the variants of coflow scheduling with rational demand matrices, for which $O(1)$-approximations were not previously known: We first present a 16-approximation algorithm for minimizing the average completion time under indirect fractional matching routings. Secondly, we address both the makespan and the average completion time objectives under indirect integral routing, and provide a full characterization of the worst-case runtimes under both objectives for this regime. While this falls short of providing $O(1)$-approximations for indirect integral routing, our proposed algorithm provides $O(\log n)$-approximations for the two objectives, in addition to being worst-case optimal; we also believe that the worst-case analysis will be helpful in determining whether a $O(1)$-approximation for indirect integral routing is possible. If it is, then we will have shown, together with the results in this paper, that constant-approximation algorithms exist for all possible coflow scheduling variants in Table˜1.